# Undergraduate data science education: Who has the microphone and what are they saying?

### v1

@@ Abstract
The presence of data science has been profound in the scientific community in almost every discipline. An important part of the data science education expansion has been at the undergraduate level. We conducted a systematic literature review to (1) specify current evidence and knowledge gaps in undergraduate data science education and (2) inform policymakers and data science educators/practitioners about the present status of data science education research. The majority of the publications in data science education that met our search criteria were available open-access. Our results indicate that data science education research lacks empirical data and reproducibility. Not all disciplines contribute equally to the field of data science education. Computer science and data science as a separate field emerge as the leading contributors to the literature. In contrast, fields such as statistics, mathematics, as well as other fields closely related to data science exhibit a limited presence in studies. We recommend that federal agencies and researchers 1) invest in empirical data science education research; 2) diversify research efforts to enrich the spectrum of types of studies; 3) encourage scholars in key data science fields that are currently underrepresented in the literature to contribute more to research and publications.
@@ 1 Introduction
The demand for data science skills in academia and industry resulted in many higher education institutions developing courses (e.g., Baumer (2015)) as well as degree programs (e.g., Glantz et al. (2023); Adhikari et al. (2021); Stern et al. (2021)). With newer educational opportunities, newer educational research questions have been formed and thus data science education emerged as a field. For instance, the American Statistical Association’s (ASA’s), Journal of Statistics Education changed its name to Journal of Statistics and Data Science Education (JSDSE) (Witmer 2020).
Over the years, a major focus of the data science education community, unsurprisingly, has been at the undergraduate level. Different professional organizations and groups have tried to describe the data science competencies including but not limited to the Park City Math Institute report (De Veaux et al. 2017), the framework of National Academies of Sciences, Engineering, and Medicine (2021), the Association for Computing Machinery’s computing competencies guidelines (Danyluk et al. 2021), the Accreditation Board for Engineering and Technology’s (ABET’s) accreditation for data science programs (2024), and the European Union’s EDISON project (Wiktorski et al. 2017). It is also worth noting that the majority of data science programs are at the master’s level and there is still a lot of room for growth at the undergraduate level (Li et al. 2021). Given the global need for data science education, it is important for us to understand undergraduate data science education as well as the scientific literature on this topic.
@@
While contributions from different educational research communities can make data science education richer, these contributions may not easily circulate across disciplines in the absence of a centralized community. For instance, Hazzan and Mike (2021) mention that despite its name JSDSE mainly caters to statistics education researchers and state that “no journal exists today that deals exclusively with data science education, let alone highlights data science education from an interdisciplinary perspective”.
The aforementioned examples of federal funding, courses, programs, curricular guidelines as well as the existence of conferences and journals with the title or keywords data science education show the growth of this field. In this manuscript, we detail a study conducted to understand data science education through in-depth readings of the existing literature. As undergraduate teacher-scholars, our focus was on undergraduate data science education.
Our goals were to (1) specify current evidence and knowledge gaps in undergraduate data science education and (2) inform policymakers and data science educators/practitioners about the present status of data science education. We conducted a systematic literature review (Evans & Benefield (2001); Liberati et al. (2009)) using criteria that we detail in Section 2. In Section 3 we share our findings and then discuss their implications in Section 4.
@@ 2 Data Collection and Analysis
We opted to extract data from six databases that potentially include publications on data science education. These databases were (1) ERIC ProQuest, (2) IEEE Xplore, (3) PubMed, (4) ScienceDirect, (5) Scopus, and (6) Web of Science. After determining the databases, we searched documents including the specific keyword “data science education” (in quotes) in at least one of the following fields: title, abstract, keywords. We found a total of 197 publications that met our criteria. Using the databases, we extracted some variables including but not limited to author names, document title, and publication venue (e.g., journal title or conference title). We conducted the initial database search in December 2022, resulting in a pool of documents that were either published by that date or available online by that date but officially published in 2023. The data on all the publications and the associated codebook for the variables are publicly available in a GitHub repository at https://github.com/mdogucu/comp-data-sci.
@@
We conducted data analysis in two stages: 1) preliminary data analysis and 2) in-depth data analysis. In the preliminary data analysis stage, we manually opened and read abstracts of publications. In the in-depth analysis stage, we read the full publications and recorded variables of potential interest. At both preliminary and in-depth analyses stages, each publication was assigned to two researchers randomly for independent review. The reviewer pairs discussed any discrepancies between their analysis decisions and tried to reach a consensus. In cases where conflicts persisted, the entire group of five researchers deliberated on the final decision.
Throughout these two stages of analysis, we excluded 41 publications due to formatting which included posters, panels, letters to journal editors, meeting highlights, publications not in English, duplicated publications, etc. We then looked at whether the remaining 156 publications were at the undergraduate level or not. For instance, if a publication focused on both K-12 and undergraduate levels, it was retained for in-depth analysis. Conversely, if a publication exclusively focused on K-12 settings, it was excluded. We excluded 18 publications that focused only on graduate level, 14 on K-12, 8 on high school, 5 on middle school and 2 on middle and high school. In total, we had 130 publications remaining by the end of the preliminary analysis and 77 publications by the end of in-depth analysis. A full list of exclusion criteria, including but not limited to the formatting and level-of-education reasons provided above, is provided in Table 1 of Supplementary Materials.
The data collection and analysis process was performed iteratively. From December 2022 to December 2023, data extracted from databases were corrected and the scope of the variables of interest was continuously revised and discussed. For instance, if a publication was not excluded based on the abstract in the preliminary stage, it could have been excluded at the in-depth analysis stage if it did not meet the inclusion criteria. Similarly, we initially extracted the open access status of publications from the databases, but when we realized many papers had multiple available versions we revised our definition of open access to include any paper accessible from a Google Scholar search.
After excluding publications on format or relevance grounds, we were left with 77 documents, which we analyzed in-depth by reading the publications fully. In addition to the variables extracted from databases (e.g. title of the publication, author names, etc.) we collected data on (1) affiliation country of researchers (2) open access status; (3) whether there were explicit research questions stated in the publication; (4) whether there was any reporting of data collection in the study and, if data were collected, the type of data (quantitative vs. qualitative); (5) content area of the publication (i.e., topic); and (6) main field and discipline of the publication. In addition, we noted big picture notes of each publication to look for patterns of main study areas of the publications.
We acknowledge that, in this study, we did not reach the entire target population of undergraduate education research publications. There may be several other publications excluded from this study that did not meet our search criteria. For instance, an article that talks about a “data science class”, “a data science activity”, or “a data science curriculum” but does not use the phrase “data science education” would be excluded from this study. A further constraint is that our target population and sample included only publications written in English. The results we share and the findings we discuss in the next two sections only pertain to the specific search criteria we used. In brief, it can be considered that we have only examined a sample of undergraduate data science education research.
@@ 3 Results
Figure 1: Undergraduate data science education publications ($n$ = 67) by year. Note that initial data retrieval was in December 2022.
Figure 1 shows the distribution of publication years for the 77 publications that we analyzed in-depth. The lone 2023 publication (Schmitt et al. 2023) was available online and obtained via our database search in December 2022, despite not being published until June 2023. The body of literature on undergraduate data science education we detected is very recent, with the oldest paper we analyzed published in 2015, and the volume of work increasing monotonically in the following years.
Over these 8 years, a majority of published work on undergraduate data science education has been freely available to the public: the breakdown of articles’ open access status is shown in Table 1 along with the document type. The vast majority of journal articles were open access, while conference publications showed a more even split. We reviewed only four publications which were neither journal nor conference articles, of which two were magazine articles (Hazzan & Mike 2021, Bonnell et al. 2022) and two were book chapters (Manzella & Emery 2022, Ryan 2016).
@@
Document Type Open Access Not Open Access Total Conference Article 25 19 44 Journal Article 25 4 29 Book Chapter 0 2 2 Magazine Article 1 1 2 Total 51 26 77
@@
We also investigated the geographic breakdown of institutions that are contributing to this body of literature. A majority of the 77 publications analyzed included at least 1 author associated with an American institution, with European and Asian institutions providing the bulk of remaining analyzed literature. Table 2 of Supplementary Materials breaks down the analyzed publications by the institutional affiliations of their authors. The higher number of scholars’ affiliations being in the United States may be at least partly attributed to the large size of the US relative to many other countries, as well as selection bias due to our search being restricted to English-written publications and to databases where US scholars might be more represented.
The majority of publications analyzed were either multidisciplinary (22 of 77 articles), treated data science as the field of study or program itself, rather than a topic in another field (18 publications), or were about computer science departments or programs (13 publications). The next most prevalent discipline was statistics, which four publications were focused on.
@@
Most publications (70 of 77 reviewed) either posed research questions and collected data or did neither, although there were exceptions. Four publications had research questions but no data collected (Vance et al. 2022, Pieterman-Bos & van Mil 2023, Hagen 2020, Robeva et al. 2020). On the other hand, three publications collected data but did not pose a specific research question or study goal (Hicks & Irizarry 2018, Rao et al. 2018, Liu & Wei 2020). Of the 38 publications that included collected data, 6, 8, and 24 publications had qualitative, quantitative, and mixed data respectively. In Table 3 of Supplementary Materials, we provide a breakdown of data collection types and the presence of research questions by discipline.
@@
Table 2: Number of articles with communicated research questions and with collected data by content area.
Research Question Collected data
Content Area
With Without Qual Quant Mixed None Total
Education Technology 6 13 1 2 5 11 19
Review Of Current State Of Data Science Education
14 2 5 1 8 2 16
Course Example 9 5 0 0 7 7 14
Pedagogical Approach
12 1 3 0 8 2 13
Program Example 4 9 1 2 0 10 13
Call To Action
1 5 0 0 0 6 6
Class Activity Example 3 3 1 1 1 3 6
Extra Curricular Activity Example
0 1 0 0 0 1 1
Guidelines 0 1 0 0 0 1 1
Note:
Some publications were classified as falling under multiple content areas.
@@
We also classified the publications based on their content area. We found publications that reviewed the current state of data science education and provided guidelines. We also encountered publications that provided program, course, class activity, and extra curricular activity examples. Some publications specifically focused on educational technology or a pedagogical approach. We also defined a content category “call to action”. While many publications had a call to action with recommendations on improving data science education, we reserved the content category “call to action” for publications that solely made a call to action, without content falling into other content areas (e.g., education technology or course example). Some publications included more than one content area and in such cases we noted down all the content areas that the publication falls into. For each content area, Table 2 shows how many publications featured that content area in a descending order, had a research question and included qualitative, quantitative or mixed type of data. In the remainder of this section, we highlight some contributions from each content area.
@@
Similar to what we are doing in our study, many scholars wanted to understand the current state of data science education. Many of the studies conducted wanted to understand and evaluate data science programs (Wiktorski et al. 2016, Oliver & McNeil 2021, Song & Zhu 2016, Shao et al. 2021, Li et al. 2021, Raj et al. 2019) or curricula (Schmitt et al. 2023). Davis (2020) specifically looked at the current state of ethics in undergraduate data science education and Ceccucci et al. (2015) considered data science education from a scientific literacy perspective. Also among these reviews were comparative studies: for instance Bile Hassan & Liu (2020)’s comparison of informatics and data science programs. Some studies also compared differences of data science approaches at country or regional levels such in Japan (Takemura 2018) or in Middle Eastern countries (Zakaria 2023). In understanding the current state of data science education, scholars also wanted to understand the data science practice. For instance, Belloum et al. (2019) developed semi-automated methods to determine the competencies needed in the job-market and Kross & Guo (2019) focused on understanding the skills that data science practitioners who also teach data science in various settings pass on to their students. In addition to understanding data science education, scholars also provided guidelines for data science degrees (Blair et al. 2021).
@@
The literature we read also included many course examples with some clever ways of including data science concepts in different courses. For instance, Fisler (2022) suggest including data science and data structures in an introductory computing course, Rao et al. (2019) teach data science through use of education data in their engineering course on modern technologies, and Haynes et al. (2019) teach data science in a general education IT course. Some institutions developed data science courses for specific disciplines including information schools (Hagen 2020) and medicine (Doudesis & Manataki 2022). Some scholars also describe courses that merge data science with philosophy of science (Pieterman-Bos & van Mil 2023) or with humanities perspectives (Vance et al. 2022). Last but not least, data science educators also tried to provide real-life data science experiences through work-integrated learning (Bilgin et al. 2022), capstone projects (Allen 2021) and entirely case-study based courses (Hicks & Irizarry 2018).
Similar to program examples and course examples provided in the literature, some publications focused on activity examples. Among class activity examples, Yamamoto et al. (2021) develop a programming exercise to bring higher-order tensor content to undergraduate level by using a 3-D puzzle. Another example is by Banadaki (2022) who develop different activities that include applications of data science in Mechanical Engineering, Biomedical science, Quantum Physics, and Cybersecurity. In data science education, there is also room for learning outside the classroom as an extra-curricular activity. This can include data hackathons (Anslow et al. 2016).
In terms of education technology, as much of data science education tackles with issues such as computing power, storage of large datasets, and sometimes automation, some of these publications focused on cloud-based data science platforms for teaching purposes (e.g., Demchenko et al. 2017). Many others focused on other tools such as online platforms, learning environments, and apps that support learning data science (e.g., Hoyt & Wangia-Anderson 2018, Bornschlegl 2016, Liu & Wei 2020, Nair et al. 2020).
@@
Almost all publications we read were calling the scientific community to action. For instance, Davis (2020) call the community to action in inclusion of ethics in undergraduate data science education but they do this through systematic review. On the other hand, we encountered publications with an important message to data science education but without a review of programs, courses, activities etc. For instance, Robeva et al. (2020) argue for the inclusion of data science in quantitative biology education. Engel (2017) draw attention to the importance of statistical literacy in data science education. We have also seen a to-do list (i.e., action list) for the community including starting a multidisciplinary data science education journal (Hazzan & Mike 2021), building a consensus on data science education and curricula (Dinov 2019), developing and deploying degree programs for university students, having basic data science training for university students, and training instructors to teach data science (Bonnell et al. 2022).

### v2

# A Systematic Literature Review of Undergraduate Data Science Education Research

@@ Abstract
The presence of data science has been profound in the scientific community in almost every discipline. An important part of the data science education expansion has been at the undergraduate level. We conducted a systematic literature review to (1) portray current evidence and knowledge gaps in self-proclaimed undergraduate data science education research and (2) inform policymakers and the data science education community about what educators may encounter when searching for literature using the general keyword ‘data science education.’ While open-access publications that target a broader audience of data science educators and include multiple examples of data science programs and courses are a strength, significant knowledge gaps remain. The undergraduate data science literature that we identified often lacks empirical data, research questions and reproducibility. Certain disciplines are less visible. We recommend that we should (1) cherish data science as an interdisciplinary field; (2) adopt a consistent set of keywords/terminology to ensure data science education literature is easily identifiable; (3) prioritize investments in empirical studies.
@@ 1 Introduction
The demand for data science skills in academia and industry resulted in many higher education institutions developing courses (e.g., Baumer 2015) as well as degree programs (e.g., Glantz et al. 2023, Adhikari et al. 2021, Stern et al. 2021). With newer educational opportunities, newer educational research questions have been formed and thus data science education emerged as a field. For instance, the American Statistical Association’s (ASA’s), Journal of Statistics Education changed its name to Journal of Statistics and Data Science Education (JSDSE) (Witmer 2020).
Over the years, a major focus of the data science education community, unsurprisingly, has been at the undergraduate level. Different professional organizations and groups have tried to describe the data science competencies including but not limited to the Park City Math Institute report (De Veaux et al. 2017), the framework of the National Academies of Sciences, Engineering, and Medicine (2018), the computing competencies guidelines of the Association for Computing Machinery’s (Danyluk et al. 2021), the accreditation for data science programs of the Accreditation Board for Engineering and Technology (ABET) (2024), and the EDISON project of the European Union (Wiktorski et al. 2017). It is also worth noting that the majority of data science programs are at the master’s level and there is still a lot of room for growth at the undergraduate level (Li et al. 2021). Given the global need for data science education, it is important for us to understand undergraduate data science education as well as the scientific literature on this topic.
@@
While contributions from different educational research communities can make data science education richer, these contributions may not easily circulate across disciplines in the absence of a centralized community. For instance, Hazzan and Mike (2021) mention that, despite its name, JSDSE mainly caters to the statistics community and state that “no journal exists today that deals exclusively with data science education, let alone highlights data science education from an interdisciplinary perspective”.
The aforementioned examples of courses, programs and curricular guidelines as well as the existence of conferences and journals with the title or keywords data science education show the growth of this field. In this manuscript, we detail a study conducted to understand undergraduate data science education through in-depth readings of the existing literature. As undergraduate teacher-scholars, our focus was on data science education at the undergraduate level. In this study, we did not intend to define data science education research or judge whether a study meets a specific definition of data science. Instead, we focused on publications that claimed to be on “data science education”.
Our goals were to (1) specify current evidence and knowledge gaps in self-proclaimed undergraduate data science education and (2) inform policymakers and the data science education community about what educators may encounter when searching for literature using the general keyword ‘data science education.’ We conducted a systematic literature review (Evans & Benefield 2001, Liberati et al. 2009) using criteria that we detail in Section 2. In Section 3 we share our findings and then we discuss their implications in Section 4.
@@ 2 Data Collection and Analysis
The target population of this literature review was publications on data science education that directly address the undergraduate level. Due to variations in terminology, keywords, and language used by teacher-scholars across different data science fields, identifying the entire target population was not feasible. Since there is no consensus on what data science is, and consequently what data science education is, we did not evaluate whether publications met a certain definition of data science. Rather, we considered publications that self-proclaim to focus on data science education. Therefore, we opted to identify the accessible population for this study as publications that included “data science education” in quotes in the title, abstract, or keywords.
Figure 1: Flowchart of data collection and analysis. Publications were collected in December 2022.
Figure 1 summarizes the stages of our data collection and analysis processes which led to the sample of this study. As shown in the diagram, we extracted data from six databases that potentially include publications on data science education. These databases were: (1) ERIC ProQuest, (2) IEEE Xplore, (3) PubMed, (4) ScienceDirect, (5) Scopus, and (6) Web of Science. These databases were selected because they cover a large number of publications and they include multiple disciplines that are commonly linked to data science education. For instance, ERIC ProQuest is known to mainly include education-focused publications, IEEE Xplore focuses on engineering and Web of Science is multidisciplinary-focused.
From the selected databases, we collected publications including the term “data science education” (in quotes) in at least one of the following fields: title, abstract, keywords. We acknowledge that there are many other combinations of search terms that could result in publications related to data science education, including terms such as “data science courses”, “data science pedagogy”, “data science curriculum”. One can generate numerous similar terms, with or without quotes. We specifically used the term “data science education” for its broadness and to set the scope of the research. We did not use “data science” and “education” separately to avoid publications in education that employ data science methods for their analyses.
@@
Across the six databases, we found a total of 197 publications that met our search criteria. We extracted some variables including, but not limited to: author names; publication title; publication venue (e.g., journal title or conference title). We conducted the initial database search in December 2022, resulting in a pool of publications that were either published by that date or available online by that date but officially published in 2023.
@@
The data analysis was conducted in two stages: 1) preliminary data analysis and 2) in-depth data analysis. Each publication was randomly assigned to two authors of the present manuscript. At both data-analysis stages, the two assigned reviewers first examined each publication independently and then discussed discrepancies between their analysis decisions, to reach a consensus. In cases where conflicts persisted, the entire group of five authors deliberated on the final decision.
During preliminary data analysis, we manually opened and read the abstracts of all publications. At this stage, we sought to exclude publications that did not meet our format and content criteria based on their abstracts. Specifically, we aimed to include journal/conference/magazine articles and book chapters (format) that focused on undergraduate data science education (content) and were written in English. During in-depth analysis we examined the full publications and, upon confirming that they met our inclusion criteria, we recorded variables of potential interest. The number of publications that were excluded due to format, content or other reason, at each stage of our analysis, is shown in Figure 1.
Across both stages, we excluded 34 publications due to formatting (which included posters, panels, letters to journal editors and meeting highlights). Of the remaining 163 publications, 77 were excluded due to their content: 12 of them were not about data science education (including e.g. publications focused on data science methodology) and 65 of them focused exclusively on a different level of education than undergraduate. Among the latter excluded group, 18 publications focused on graduate level, 29 focused on K-12, middle school or high school and 18 publications focused on data science education for adults in non-academic programs (including, e.g. practitioners, citizen science and instructors). Publications that focused on both undergraduate and non-undergraduate levels were included. Of the remaining 86 publications, other 9 were excluded due to being: not written in English; duplicated in our dataset; retracted by their authors.
After excluding publications for the reasons detailed above, we were left with 77 publications, which we analyzed in-depth. In addition to the variables extracted from databases (e.g. title of the publication, author names, etc.) for each publication we collected data on:
  • affiliation country of researchers
  • open access status (i.e. whether the full publication is accessible for free from a Google Scholar search)
  • year when the publication was first published online
  • document type (conference article, journal article, magazine article or book chapter)
  • whether there were explicit research questions stated in the publication
  • whether there was any reporting of data collection in the publication and, if data were collected, the type of data (quantitative, qualitative or mixed)
  • publication focus, that is a categorization of the subject matter of the publication (for example “pedagogical approach”, “class activity” or “review of current state of data science education”)
  • discipline of the publishing source, determined by examining the call for contributions of the journal or conference, or the description of the book or magazine (“broad” when the publishing source called for contributions across all data science fields, otherwise a specific sub-field of data science, e.g. “computer science” or “statistics”)
  • the discipline of the target audience, as expressed by the authors in the publication (“broad” when the publication target were all data science educators, otherwise a specific sub-field of data science, e.g. “computer science” or “statistics”)
@@ 3 Results
Figure 2: Undergraduate data science education publications ($n$ = 77) by year they were first made available online. Note that initial data retrieval was in December 2022.
Publication years. Figure 2 shows the distribution of the years when publications in our review were made available online. Some of these publications were available during our database search in December 2022, but would ultimately be published in a 2023 journal edition (e.g., Pieterman-Bos & van Mil 2023, Schmitt et al. 2023). The body of literature that we identified is very recent, with the oldest paper available in 2015 and the volume of work increasing steadily in the following years.
Open access status and document types. A majority of work on undergraduate data science education published over these 8 years has been freely available to the public: the breakdown of articles’ open access status is shown in Table LABEL:tbl-open_access along with the document type. The vast majority of journal articles were open access, while conference publications showed a more even split. We reviewed only four publications which were neither journal nor conference articles, of which two were magazine articles (Hazzan & Mike 2021, Bonnell et al. 2022) and two were book chapters (Manzella & Emery 2022, Ryan 2016).
@@
Document Type Open Access Not Open Access Total
Conference Article 24 19 43 (56%)
Journal Article 26 4 30 (39%)
Book Chapter 0 2 2 (3%)
Magazine Article 1 1 2 (3%)
Total 51 (66%) 26 (34%) 77
@@
Affiliation countries. We also investigated the geographic breakdown of institutions that are contributing to literature on undergraduate data science education. A majority of the publications analyzed included at least 1 author associated with an American institution (45 of 77), with European and Asian institutions providing the bulk of the remaining analyzed literature. Table LABEL:tbl-country-affiliation breaks down the analyzed publications by the institutional affiliations of their authors. The higher number of scholars’ affiliations being in the United States may be at least partly attributed to the large size of the US relative to many other countries, as well as selection bias due to our search being restricted to English-written publications and to databases where US scholars might be more represented.
Table 2: For each country below, we report the number of publications which included at least one affiliated author. Affiliation country of authors was determined based on their home institutions as reported in the publication, and does not represent authors’ nationalities.
Country # Publications Country # Publications
United States of America 45 Australia 2
Netherlands 8 Egypt 2
Canada 5 Germany 2
Norway 5 Greece 1
United Kingdom 5 Hong Kong S.A.R. 1
Spain 4 Ireland 1
China 3 New Zealand 1
India 3 Portugal 1
Israel 3 Romania 1
Italy 3 Slovakia 1
Japan 3 Thailand 1
Switzerland 3 United Arab Emirates 1
Figure 3: Composition of the discipline of targeted audiences of publications on data science education (right) by the discipline of the publishing source (left). A total of 2 publications were not included in the plot, one because it stated to target both statistics and computer science, the other one because it was not possible to determine its audience discipline.
Source and audience disciplines. Figure 3 shows the breakdown of articles by discipline. We report both the discipline of the publishing venue (such as the journal’s or conference’s discipline), and the audience that the publication stated (or implied) to target. About half of the articles (40/77) were written for the broad community of data science educators. The remaining articles were written for educators in particular disciplines, most commonly computer science (12 publications), statistics (5 publications), and information sciences (5 publications). We found that publication venues in all of these disciplines, as well as engineering, contributed articles aimed at their specific audience as well as articles aimed at the broad community of data science educators.
@@
Research Question and Data Collection. Most publications (69 of 77 reviewed) either posed research questions and collected data or did neither, although there were exceptions. Four publications had research questions but no data collected (Vance et al. 2022, Pieterman-Bos & van Mil 2023, Hagen 2020, Robeva et al. 2020). On the other hand, four publications collected data but did not pose a specific research question or study goal (Hicks & Irizarry 2018, Rao et al. 2018, Liu & Wei 2020, Cuadrado-Gallego et al. 2021). Of the 40 publications that included collected data, 6, 9, and 25 publications had qualitative, quantitative, and mixed data respectively.
@@
Figure 4: Number of publications with communicated research questions and with collected data stratified by publication focus.
@@
Publication focus. We also classified publications based on their focus. We found publications that reviewed the current state of data science education and provided guidelines. We also encountered publications that provided examples of programs, courses, class activities, and extracurricular activities. Some publications specifically focused on educational technology or discussed a pedagogical approach. Some publications included more than one focus: in such cases, we noted down all the categories that the publication falls into, with the exception of call to action. Almost all publications we read were calling the scientific community to action and this is unsurprising for scientific publications. For instance, an educational technology paper might call to action to use a specific tool and a pedagogy paper might call to action to use a specific pedagogical approach. However, some publications solely made a call to action, without content falling into other publication focuses (e.g., education technology or course example). For these publications, we reserved the publication focus “call to action”, that is, papers that we labeled in the “call to action” category did not fall in any other category for publication focus.
For each category of publication focus, Figure 4 shows how many publications featured that publication focus, had a research question and included qualitative, quantitative or mixed type of data. Note that, in this figure, publication focuses are arranged and grouped to match the order in which we are going to illustrate them below, as we highlight contributions from each publication focus.
@@
Similar to what we are doing in our study, many scholars wanted to understand the current state of data science education. Many of the studies conducted wanted to understand and evaluate data science programs (Wiktorski et al. 2016, Oliver & McNeil 2021, Song & Zhu 2016, Shao et al. 2021, Li et al. 2021, Raj et al. 2019) or curricula (Schmitt et al. 2023). Davis (2020) specifically looked at the current state of ethics in undergraduate data science education and Ceccucci et al. (2015) considered data science education from a scientific literacy perspective. Among these reviews, there were also comparative studies: for instance, Bile Hassan & Liu (2020)’s comparison of informatics and data science programs. Some studies also compared differences of data science approaches at country or regional levels such as in Japan (Takemura 2018), or in Middle Eastern countries (Zakaria 2023). In understanding the current state of data science education, scholars also wanted to understand the data science practice. For example, Belloum et al. (2019) developed semi-automated methods to determine the competencies needed in the job-market and Kross & Guo (2019) focused on understanding the skills that data science practitioners who teach data science in various settings pass onto their students. In addition to reviewing data science education, scholars also provided guidelines for data science degrees (Blair et al. 2021).
@@
The literature we read also included many course examples with some clever ways of including data science concepts in different courses. For instance, Fisler (2022) suggested to include data science and data structures in an introductory computing course, Rao et al. (2019) teach data science through the use of education data in their engineering course on modern technologies, and Haynes et al. (2019) teach data science in a general education IT course. Some institutions developed data science courses for specific disciplines including information schools (Hagen 2020) and medicine (Doudesis & Manataki 2022). Some scholars also described courses that merge data science with philosophy of science (Pieterman-Bos & van Mil 2023) or with humanities perspectives (Vance et al. 2022). Last but not least, data science educators also tried to provide real-life data science experiences through work-integrated learning (Bilgin et al. 2022), capstone projects (Allen 2021) and entirely case-study based courses (Hicks & Irizarry 2018).
Similar to program examples and course examples provided in the literature, some publications focused on activity examples. Among class activity examples, Yamamoto et al. (2021) developed a programming exercise to bring higher-order tensor content to undergraduate level by using a 3-D puzzle. Another example is by Banadaki (2022) who developed different activities that include applications of data science in Mechanical Engineering, Biomedical science, Quantum Physics, and Cybersecurity. In data science education, there is also room for learning outside the classroom as an extra-curricular activity. This included data hackathons (Anslow et al. 2016).
In terms of education technology, as much of data science education tackles with issues such as computing power, storage of large datasets, and sometimes automation, some of these publications focused on cloud-based data science platforms for teaching purposes (e.g., Demchenko et al. 2017). Many others focused on tools such as online platforms, learning environments, and apps that support learning data science (e.g., Hoyt & Wangia-Anderson 2018, Bornschlegl 2016, Liu & Wei 2020, Nair et al. 2020).
@@
Almost all publications we read were calling the scientific community to action. For instance, through a systematic review, Davis (2020) called the community to action to include ethics in undergraduate data science education. We also encountered publications with an important message about data science education but without a review of programs, courses, activities etc. For instance, Robeva et al. (2020) argued for the inclusion of data science in quantitative biology education. Engel (2017) drew attention to the importance of statistical literacy in data science education. We have also seen a to-do list (i.e., action list) for the community including starting a multidisciplinary data science education journal (Hazzan & Mike 2021), building a consensus on data science education and curricula (Dinov 2019), developing and deploying degree programs for university students, having basic data science training for university students, and training instructors to teach data science (Bonnell et al. 2022).